OpenStudy (anonymous):

PLEASE HELP! Prove: tan2 θ•cos2 θ + cos2 θ = 1. You must show all work

OpenStudy (zzr0ck3r):

hint: \(\tan^2(\theta)\cos^2(\theta)=\frac{\sin^2(\theta)}{\cos^2(\theta)}\cos^2(\theta)=\frac{\sin^2(\theta)}{\cancel{\cos^2(\theta)}}\cancel{\cos^2(\theta)}\)

OpenStudy (anonymous):

Is sin^2 T = 1??

OpenStudy (zzr0ck3r):

another hint: There was a dude named Pythagoras, and he and his friends did some pretty cool stuff.

OpenStudy (anonymous):

Haha I understand, what do I put into the pythagorean thereom?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

So sin^2 T = 1 is the final portion of this answer?

OpenStudy (zzr0ck3r):

\(\tan^2(\theta)\cos^2(\theta)+\cos^2(\theta)=\frac{\sin^2(\theta)}{\cos^2(\theta)}\cos^2(\theta)+\cos^2(\theta)=\\\frac{\sin^2(\theta)}{\cancel{\cos^2(\theta)}}\cancel{\cos^2(\theta)}+\cos^2(\theta)=\sin^2(\theta)+\cos^2(\theta)= \ \huge ?\)

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

correct, that is the Pythagorean theorem \(\cos^2(x)+\sin^2(x) = 1\) for all \(x\).

OpenStudy (anonymous):

Thank you so much @zzr0ck3r

OpenStudy (zzr0ck3r):

np.